# 11footnotetext: Corresponding authors22footnotetext: Project lead

### v1

@@ Appendix E Additional ablation experiments
Reasons behind channel-wise control in VMC. In the main paper, we study how to control the auxiliary matrix to produce a compatible $\sigma_{\mathbf{R}}$ by searching for the $r^{\alpha}$. As for the variance ratio $r$, it is worth noting that we adopt a per-channel scaling factor $r\in\mathbf{R}^{C_{out}}$ instead of the per-tensor one. Here we elaborate on the reasons for this. As shown in Fig. 11, we give a boxplot of the pre-trained weight $\mathbf{W}$ and it can be seen that the values for each channel exhibit different variances. Therefore, considering that the auxiliary matrix has a large impact on performance, we opt for this fine-grained channel-wise control. In addition, since $r$ is task-agnostic, it can be computed in advance and stored once. Finally, we give the distribution of $r$, and one can see that $r$ is generally less than 1, with a single peak and a left-skewed distribution shape.
Figure 11: The value ranges of different channels in the pre-trained weights.
Appendix F Advantage Discussion of IntLoRA.
Here, we would like to discuss the advantages of our InLoRA, to highlight the contribution of this work. First, since the downstream fine-tuning is performed on the quantized pre-trained weights, and thus the adaptation pipeline with our IntLoRA only needs to load the quantized weights, reducing memory footprint from FP weights. Second, transferring the adaptation parameters to also the INT arithmetic allows all weights of the model can be stored in INT format, reducing the disk space for storage. Thirdly, thanks to the designed multiplicative adaptation form, our IntLoRA enables seamless weight merge of quantized pre-trained weights and adaptation weights at test time without additional post-training quantization, streamlining the adaptation process while alleviating performance drop from quantization.
@@ Appendix G Limitation and future works
Although the proposed IntLoRA can effectively improve the efficiency of diffusion model fine-tuning by allowing the adaptation parameters also on the integer arithmetic, the proposed framework can be further improved in the following aspects. First, although the trainable low-rank weights are quantized with STE, these quantized weights are still in FP type during tuning to enable accurate gradient updates. Therefore, it is promising to specifically design integer-type propagation. Despite this seems challenging, it can further reduce the training cost and accelerate the adaptation process. Second, although we introduce a feasible way that uses hyperparameter search of the smoothing factor $\alpha$ to find a compatible $\sigma_{\mathbf{R}}$ as well as the appropriate distribution shape of $\mathbf{R}$, it can be more elegant if we can use advanced mathematical analysis techniques, such as functional analysis, to find the statistical properties a suitable $\mathbf{R}$ should satisfy. Third, this work mainly focuses on the efficient acceleration of LoRA due to its prevalence among the PEFT techniques. Applications to other PEFT methods for hardware-efficient adaptation could be interesting future work.
@@ Appendix H Additional visualization results
Figure 18: The style images used for training and the text prompts for evaluation.

### v3

@@ Appendix E Impacts from the Auxiliary Matrix
Appendix F Justification for the Value Orders
A key assumption in the derivation for VMC is that the learned values of low-rank parameters $\mathbf{AB}$ are orders of magnitude smaller than the auxiliary matrix $\mathbf{R}$. Based on this assumption we ignore $\mathbf{AB}$ as an approximation. Here, we give the specific evidence for this approximation. Specifically, we visualize the weights of the trained $\mathbf{AB}$ and the distribution of $\mathbf{R}$, as shown in Fig. 12. It can be seen the range of $\mathbf{AB}$ is constrained to $[-0.0004,0.0004]$, while the range of $\mathbf{R}$ is $[-0.08,008]$. Therefore, the experimental visualization above confirms the soundness of our approximation. Since the $\mathbf{AB}$ in LoRA is zero-initialized, it tends to be distributed around zero with learned small values aiming to not disturb the pre-training weights too much.
@@ Appendix G Limitation and Future Works
Although the proposed IntLoRA can effectively improve the efficiency of diffusion model fine-tuning by allowing the adaptation parameters also on the integer arithmetic, the proposed framework can be further improved in the following aspects. First, although the trainable low-rank weights are quantized with STE, these quantized weights are still in FP16 type during tuning to enable accurate gradient updates. Therefore, it is promising to specifically design integer-type propagation. Despite this seems challenging, it can further reduce the training cost and accelerate the adaptation process. Second, although we introduce a feasible way that uses hyperparameter search of the smoothing factor $\alpha$ to find a compatible $\sigma_{\mathbf{R}}$ as well as the appropriate distribution shape of $\mathbf{R}$, it can be more elegant if we can use advanced mathematical analysis techniques, such as functional analysis, to find the statistical properties a suitable $\mathbf{R}$ should satisfy. Third, this work mainly focuses on the efficient acceleration of LoRA due to its prevalence among the PEFT techniques. Applications to other PEFT methods for hardware-efficient adaptation could be interesting future work.
@@ Appendix H Additional Visualization Results
Figure 19: The style images used for training on the style-customized generation task. The text prompts are “A friendly robot in [V] style”, “A fluffy baby sloth with a knitted hat trying to figure out a laptop in [V] style”, “A panda eating bamboo in [V] style”, “The Golden Gate bridge in [V] style”, “A man riding a snowboard in [V] style”, “The letter ‘G’ in [V] style”.